# Solving dynamic portfolio selection problems via score-based diffusion models

### v1

@@ 1.3 Notations
The rest of this paper is organized as follows. In Section 2 we present a adaptive training and sampling framework of score-based diffusion models, and provide a $\mathcal{A}\mathcal{W}_{2}$-quantification error bound between the output distribution and data distribution. In Section 3, we turn to a specific mean-variance portfolio selection problem, and establish the model stability in terms of the adapted Wasserstein metric (i.e., the value function is Lipschitz in the adapted Wasserstein metric). Section 4 focuses on applications of the theoretical results and in particular, develops a policy gradient algorithm based on generative environment along with experimental results. 5 concludes the main context. Appendices include technical discussions and proofs of the main results in Section 2.
@@ 2.1 Conditional score matching
For simplicity, we take OU process as the forward process of our diffusion model:
@@
We remark that (2.3) is the classical score error bound for the first marginal of $\mathbb{P}$; see Song et al. (2021b). (2.4) is crucial to our adaptive sampling scheme because it provides a proxy to $\nabla_{x}\log p_{t+1}(\tau,\cdot|X^{1:t})$, the score function conditioning on the realization $X^{1:t}$. (2.4) is feasible due to the following denoising score-matching techniques. The proof is somewhat similar to classical denoising score matching (e.g., Hyvärinen (2005)), yet here we consider a conditional version. We provide a proof for self-containedness.
@@
$-2\mathbb{E}_{X^{1:t}\sim\mathbb{P}_{1:t}}\int_{\mathbb{R}^{d}}s^{t+1}_{\theta}(\tau,X^{1:t},x)\cdot\nabla_{x}\int_{\mathbb{R}^{d}}p(\tau,x|x_{0})\mathbb{P}_{X^{1:t}}(\mathrm{d}x_{0})\mathrm{d}x$ (2.15)
@@
$-2\mathbb{E}_{X^{1:t}\sim\mathbb{P}_{1:t}}\mathbb{E}_{X^{t+1}\sim\mathbb{P}_{X^{1:t}}}\mathbb{E}_{X_{\tau}^{t+1}\sim p(\tau,\cdot|X^{t+1})}\big{[}s^{t+1}_{\theta}(\tau,X^{1:t},X_{\tau}^{t+1})\cdot\nabla_{x}p(\tau,X_{\tau}^{t+1}|X^{t+1})\big{]}$ (2.17)
@@
$-2\mathbb{E}_{X^{1:t+1}\sim\mathbb{P}_{1:t+1}}\mathbb{E}_{X_{\tau}^{t+1}\sim p(\tau,\cdot|X^{t+1})}\big{[}s^{t+1}_{\theta}(\tau,X^{1:t},X_{\tau}^{t+1})\cdot\nabla_{x}p(\tau,X_{\tau}^{t+1}|X^{t+1})\big{]}$ (2.19)
$=$ $\mathbb{E}_{X^{1:t+1}\sim\mathbb{P}_{1:t+1}}\mathbb{E}_{X_{\tau}^{t+1}\sim p(\tau,\cdot|X^{t+1})}\bigg{|}s^{t+1}_{\theta}(\tau,X^{1:t},X_{\tau}^{t+1})-\nabla_{x}\log p(\tau,X_{\tau}^{t+1}|X^{t+1})\bigg{|}^{2}+C.$ (2.20)
@@ 2.2 An adaptive sampling method
Suppose the real model of time series data is $\mathbb{P}$, and we have a pre-trained diffusion model $s_{\theta}^{t}$ satisfies (2.3) and (2.4). We note that, fixing $x^{1:t}$, $s^{t+1}_{\theta}(\tau,x^{1:t},\cdot)$, serves as an approximation of the conditional score function $\nabla_{x}\log p(\tau,\cdot|x^{1:t})$. This motives us to sample from the output distribution adaptively; see Algorithm 1.
@@
1:  for $n=1$ to $N$ do
2:      With initial condition $Y_{0}=Z_{(n)}^{1}$, run the backward SDE with score matching function $s^{\theta}_{1}$.
3:      $y_{(n)}^{1}\leftarrow Y_{\mathcal{T}}$.
4:     for $t=1$ to $T-1$ do
5:         With initial condition $Y_{0}=Z_{(n)}^{t+1}$, run the backward SDE with score matching function $(\tau,x)\mapsto s^{t+1}_{\theta}(\tau,y_{(n)}^{1:t},x)$.
6:         $y_{(n)}^{t+1}\leftarrow Y_{\mathcal{T}}$.
7:     end for
@@
We now proceed to prove that the sampling method described by Algorithm 1 gives a distribution $\mathbb{Q}$ close to $\mathbb{P}$ in $\mathcal{A}\mathcal{W}_{2}$. Our proof relies on the existing total variation bounds of diffusion models between conditional distributions and an explicit tail estimate of SDEs. We need the following technical assumptions for our results.
@@
    There exists constants $C,R_{0},\delta>0$ such that for any for any $\theta\in\Theta$, $t\in\{1,2,\cdots,T\}$, $\tau\in[0,\mathcal{T}]$, $x^{1:t}\in\mathbb{R}^{dt}$ and $x\in\mathbb{R}^{d}$ with $|x|>R_{0}$,
    $2x\cdot s^{1}_{\theta}(\tau,x)\leq-(1+\delta)|x|^{2}+C,$ (2.27)
    $2x\cdot s^{t+1}_{\theta}(\tau,x^{1:t},x)\leq-(1+\delta)|x|^{2}+C.$ (2.28)
Remark 2.
Typically, it is advisable to refrain from making assumptions about the approximation network, as such assumptions may potentially compromise its approximating power. However, in this paper, we demonstrate that by adhering to Assumption 2, which are all assumptions pertaining to the data distribution, we can select the approximating network $s_{\theta}$ such that it satisfies Assumption 3 while maintaining the approximation errors (2.3) and (2.4). To avoid deviating from the primary objective of this paper, we present these results in Appendix A.
@@
For $t\geq 1$, denote by $\{X^{x^{1:t}}_{\tau}\}_{\tau\in[0,\mathcal{T}]}$ the forward process with initial condition $X^{x^{1:t}}_{0}\sim\mathbb{P}_{x^{1:t}}$, and $\{Y^{y^{1:t}}_{\tau}\}_{\tau\in[0,\mathcal{T}]}$ the backward process with initial condition $Y^{y^{1:t}}_{0}\sim\mathcal{N}(0_{\mathbb{R}^{d}},I_{d\times d})$ and score matching function $(\tau,x)\mapsto s^{t}_{\theta}(\tau,y^{1:t},x)$. Define $\mathbb{Q}^{\mathcal{T}}_{y^{1:t}}=\mathcal{L}(Y^{y^{1:t}}_{\mathcal{T}})$. Moreover, define $\mathbb{Q}^{\mathcal{T}}_{1}$ to be output distribution of backward process with score matching function $s_{\theta}^{1}$.
@@
    $\alpha(\mathcal{T}):=\mathcal{T}^{2}e^{-\mathcal{T}}+e^{-c\mathcal{T}}.$ (2.31)
@@
    $\mathbb{Q}^{\mathcal{T}}(\mathrm{d}x_{1},\mathrm{d}x_{2},\cdots,\mathrm{d}x_{T}):=\mathbb{Q}^{\mathcal{T}}_{1}(\mathrm{d}x_{1})\mathbb{Q}^{\mathcal{T}}_{x_{1}}(\mathrm{d}x_{2})\cdots\mathbb{Q}^{\mathcal{T}}_{x^{1:t}}(\mathrm{d}X^{t+1})\cdots\mathbb{Q}^{\mathcal{T}}_{x^{1:T-1}}(\mathrm{d}x_{T}).$ (2.32)
@@
    Because $\alpha(\mathcal{T})$ decays exponentially when $\mathcal{T}\to\infty$, we have
    $\mathcal{T}^{\frac{5(T-1)}{2}}\alpha(\mathcal{T})^{1/2^{T-1}}\to 0,\quad\mathcal{T}\to\infty.$
    In other words, Equation (2.33) offers a convergence guarantee provided that $\mathcal{T}$ is sufficiently large and $\varepsilon_{\rm score}$ is sufficiently small. Even in the static data scenario, achieving convergence under the Wasserstein metric for the diffusion model is recognized as challenging; see Remark 1 of Kwon et al. (2022). Under various technical assumptions on the data distribution, such as strong log-concavity, there are Wasserstein convergence results available (e.g., Gao et al. (2025), Tang and Zhao (2024a)), and it is possible to bound $\mathcal{A}\mathcal{W}_{2}$ by $\mathcal{W}_{2}$ up to some explicit error terms; see Blanchet et al. (2024). However, with this conventional approach, obtaining bounds for conditional distributions requires imposing strong log-concavity on every conditional distribution of $\mathbb{P}$. We circumvent these assumptions and replace them with certain disspative conditions of the networks, which are attained by carefully designing the network structure (i.e., Assumptions 3).
@@
    It is evident that, as outlined in Algorithm 1, $\{y_{(n)}^{1}\}_{n=1}^{N}$ is sampled from the distribution $\mathbb{Q}_{1}$. Furthermore, given the sequence $y_{(n)}^{1:t}$, the next sample $y_{(n)}^{t+1}$ is drawn from the distribution $\mathbb{Q}_{y_{(n)}^{1:t}}$. By definition, $\{y^{(n)}_{1},y^{(n)}_{2},\cdots,y^{(n)}_{T}\}_{n=1}^{N}$ is one sample from the distribution $\mathbb{Q}$, thereby establishing that $\mathbb{Q}$ is the output distribution of Algorithm 1. Consequently, Theorem 2.2 provides $\mathcal{A}\mathcal{W}_{2}$ bounds between the output distribution of Algorithm 1 and the data distribution.
@@
    Due to the construction of $\mathbb{Q}$ (as defined in (2.32)), if $Y^{1:t}\sim\mathbb{Q}$, $\mathbb{Q}_{y^{1:t}}$ is the conditional distribution of $Y^{t+1}$ given $Y^{1:t}=y^{1:t}$. Consequently, it can be sampled by selecting the score matching function $s^{t+1}_{\theta}$. In essence, Algorithm 1 can also be utilized for conditional sampling, provided that $y^{1:t}$ is included as input, and the sampling process starts with $y^{(n)}_{t+1}$.
@@
Consequently, for any $\varepsilon,p>0$, if $\mathcal{T}$ is sufficiently large such that $\alpha(\mathcal{T})<\varepsilon p/(2C)$, and $\varepsilon_{\rm score}<\frac{\varepsilon p\mathcal{T}^{-\frac{5}{2}}}{2C}$, then with probability at least $1-p$, we have
@@
provided that the prior $X^{1:t}$ is randomly sampled from $\mathbb{P}_{X^{1:t}}$.

### v3

@@ 1.4 Frequently used notation
We use $K,C,c$ to denote generic constants which may vary from line to line.
@@ 2.1 Conditional score matching
For simplicity, we take the OU process as the forward process of our diffusion model:
@@
We remark that (2.3) is the classical score error bound for the first marginal of $\mathbb{P}$; see Song et al. (2021b). (2.4) is crucial to our adaptive sampling scheme because it provides a proxy to $\nabla_{x}\log p_{t+1}(\tau,\cdot|X^{1:t})$, the score function conditioning on the realization $X^{1:t}$. However, (2.4) is not feasible directly, because we can not sample from the conditional probability, but rather, only have access to the joint distribution (data distribution). This issue can be resolved by the following denoising score-matching techniques, in which the error can be computed using only joint distribution. While the idea is borrowed from classical denoising score matching (e.g., Hyvärinen (2005)), we need to consider a conditional version and the proof is more complicated.
@@
$-2\mathbb{E}_{X^{1:t}\sim\mathbb{P}_{1:t}}\int_{\mathbb{R}^{d}}s^{t+1}_{\theta}(\tau,X^{1:t},x)\cdot\nabla_{x}\int_{\mathbb{R}^{d}}\phi(\tau,x|x_{0})\mathbb{P}_{X^{1:t}}(\mathrm{d}x_{0})\mathrm{d}x$ (2.15)
@@
$-2\mathbb{E}_{X^{1:t}\sim\mathbb{P}_{1:t}}\mathbb{E}_{X^{t+1}\sim\mathbb{P}_{X^{1:t}}}\mathbb{E}_{X_{\tau}^{t+1}\sim\phi(\tau,\cdot|X^{t+1})}\big{[}s^{t+1}_{\theta}(\tau,X^{1:t},X_{\tau}^{t+1})\cdot\nabla_{x}\phi(\tau,X_{\tau}^{t+1}|X^{t+1})\big{]}$ (2.17)
@@
$-2\mathbb{E}_{X^{1:t+1}\sim\mathbb{P}_{1:t+1}}\mathbb{E}_{X_{\tau}^{t+1}\sim\phi(\tau,\cdot|X^{t+1})}\big{[}s^{t+1}_{\theta}(\tau,X^{1:t},X_{\tau}^{t+1})\cdot\nabla_{x}\phi(\tau,X_{\tau}^{t+1}|X^{t+1})\big{]}$ (2.19)
$=$ $\mathbb{E}_{X^{1:t+1}\sim\mathbb{P}_{1:t+1}}\mathbb{E}_{X_{\tau}^{t+1}\sim\phi(\tau,\cdot|X^{t+1})}\bigg{|}s^{t+1}_{\theta}(\tau,X^{1:t},X_{\tau}^{t+1})-\nabla_{x}\log\phi(\tau,X_{\tau}^{t+1}|X^{t+1})\bigg{|}^{2}+C.$ (2.20)
@@ 2.2 An adaptive sampling method
Suppose the real model of time-series data is $\mathbb{P}$, and we have a pre-trained diffusion model $s_{\theta}^{t}$ satisfies (2.3) and (2.4). We note that, fixing $x^{1:t}$, $s^{t+1}_{\theta}(\tau,x^{1:t},\cdot)$, serves as an approximation of the conditional score function $\nabla_{x}\log p(\tau,\cdot|x^{1:t})$. This motives us to sample from the output distribution adaptively; see Algorithm 1.
@@
1:for $n=1$ to NN do
2:   With initial condition $\bar{Y}_{0}=Z_{(n)}^{1}$, run the backward SDE with score matching function $s^{\theta}_{1}$. , i.e.,
$\mathrm{d}\bar{Y}_{\tau}=[\bar{Y}_{\tau}+2s_{\theta}^{1}(\mathcal{T}-\tau,\bar{Y}_{\tau})]\mathrm{d}\tau+\sqrt{2}\bar{B}_{\tau}.$ (2.22)
3:   $y_{(n)}^{1}\leftarrow\bar{Y}_{\mathcal{T}}$.
4:  for $t=1$ to T1T-1 do
5:   $y^{1:t}\leftarrow y^{1:t}_{(n)}$
6:    With initial condition $Y_{0}=Z_{(n)}^{t+1}$, run the backward SDE with score matching function $(\tau,x)\mapsto s^{t+1}_{\theta}(\tau,y^{1:t},x)$. , i.e,
$\mathrm{d}\bar{Y}_{\tau}=[\bar{Y}_{\tau}+2s_{\theta}^{t+1}(\mathcal{T}-\tau,y^{1:t},\bar{Y}_{\tau})]\mathrm{d}\tau+\sqrt{2}\bar{B}_{\tau}.$ (2.23)
7:    $y_{(n)}^{t+1}\leftarrow\bar{Y}_{\mathcal{T}}$.
@@
We now proceed to prove that the sampling method described by Algorithm 1 gives a distribution $\mathbb{Q}$ close to $\mathbb{P}$ in $\mathcal{A}\mathcal{W}_{2}$. Our proof relies on the existing total variation bounds of diffusion models between conditional distributions and an uniform tail estimate of SDEs. We need the following technical assumptions for our results.
@@
    There exist constants $M_{\rm disp},\delta>0$ such that for any for any $t\in\{1,2,\cdots,T\}$, $\tau\in[0,\mathcal{T}]$, $x^{1:t}\in\mathbb{R}^{dt}$ and $x\in\mathbb{R}^{d}$,
    $2x\cdot s^{1}_{\theta}(\tau,x)\leq-(1+\delta)|x|^{2}+{M_{\rm disp}},$ (2.29)
    $2x\cdot s^{t+1}_{\theta}(\tau,x^{1:t},x)\leq-(1+\delta)|x|^{2}+{M_{\rm disp}}.$ (2.30)
Typically, it is advisable to refrain from making assumptions about the approximation network, as such assumptions may potentially compromise its approximating power. However, in this paper, we demonstrate that by adhering to Assumption 2, which are all assumptions pertaining to the data distribution, we can select the approximating network $s_{\theta}$ such that it satisfies Assumption 3 (with appropriately chosen $M_{\rm disp}$), while maintaining the approximation errors (2.3) and (2.4). See Proposition 2.2 below. To avoid deviating from the primary objective of this paper, we present detailed discussions on this result in Appendix A.
Proposition 2.2.
Suppose Assumption 2 holds. For any $\varepsilon_{\rm score}>0$, there exists a $s_{\theta}$ satisfying Assumption 1 as well as Assumption 3 with $M_{\rm disp}\sim\log(1/\varepsilon_{\rm score})$.
We are now ready to present one of our main results: the quantitative error bounds between output distribution of the adaptive sampling method, and the real data distribution.
@@
For $t\geq 1$, denote by $\{X^{x^{1:t}}_{\tau}\}_{\tau\in[0,\mathcal{T}]}$ the forward process with initial condition $X^{x^{1:t}}_{0}\sim\mathbb{P}_{x^{1:t}}$, and $\{Y^{y^{1:t}}_{\tau}\}_{\tau\in[0,\mathcal{T}]}$ the backward process with initial condition $Y^{y^{1:t}}_{0}\sim\mathcal{N}(0_{\mathbb{R}^{d}},I_{d\times d})$ and score matching function $(\tau,x)\mapsto s^{t}_{\theta}(\tau,y^{1:t},x)$, i.e., $Y^{y^{1:t}}$ satisfies (2.23). Define $\mathbb{Q}^{\mathcal{T}}_{y^{1:t}}=\mathcal{L}(Y^{y^{1:t}}_{\mathcal{T}})$. Moreover, define $\mathbb{Q}^{\mathcal{T}}_{1}=\mathcal{L}(Y_{\mathcal{T}})$ with $Y$ satisfying (2.22).
@@
    $\alpha(\mathcal{T}):=\mathcal{T}^{2}e^{-\mathcal{T}}+e^{-c\mathcal{T}+cM_{\rm disp}}.$ (2.33)
@@
    $\mathbb{Q}^{\mathcal{T}}(\mathrm{d}x^{1},\mathrm{d}x^{2},\cdots,\mathrm{d}x^{T}):=\mathbb{Q}^{\mathcal{T}}_{1}(\mathrm{d}x^{1})\mathbb{Q}^{\mathcal{T}}_{x^{1}}(\mathrm{d}x^{2})\cdots\mathbb{Q}^{\mathcal{T}}_{x^{1:t}}(\mathrm{d}x^{t+1})\cdots\mathbb{Q}^{\mathcal{T}}_{x^{1:T-1}}(\mathrm{d}x^{T}).$ (2.34)
@@
    By Proposition 2.2, with vanishing $\varepsilon_{\rm score}$ we can take $M_{\rm disp}\sim\log(1/\varepsilon_{\rm score})$, thus $\alpha(\mathcal{T})\leq\mathcal{T}^{2}e^{-c\mathcal{T}+cM_{\rm disp}}\leq\mathcal{T}^{2}\varepsilon_{\rm score}^{-1}e^{-c\mathcal{T}}$. In this case (2.35) reads
    $\mathcal{A}\mathcal{W}_{2}^{2}(\mathbb{P},\mathbb{Q}^{\mathcal{T}})\leq C\Big{(}\mathcal{T}^{\frac{5T}{2}}\varepsilon_{\rm score}^{1/2^{T-1}}+\mathcal{T}^{\frac{5(T-1)}{2}+\frac{1}{2^{T-2}}}e^{-c\mathcal{T}/2^{T-1}}\varepsilon_{\rm score}^{-1/2^{T-1}}\Big{)}.$ (2.36)
    Thus, taking $\varepsilon_{\rm score}=\mathcal{T}^{-52^{T-3}+1}e^{-c\mathcal{T}/2}$, (2.36) becomes
    $\mathcal{A}\mathcal{W}^{2}_{2}(\mathbb{P},\mathbb{Q}^{\mathcal{T}})\leq C\mathcal{T}^{\frac{5T}{2}-\frac{5}{4}+\frac{1}{2^{T-1}}}e^{-\frac{c\mathcal{T}}{2^{T}}}.$ (2.37)
    which provides a convergence guarantee when $\mathcal{T}\to\infty$ (equivalently, $\varepsilon_{\rm score}\to 0$). Even in the static data scenario, achieving convergence under the Wasserstein metric for the diffusion model is recognized as challenging; see Remark 1 of Kwon et al. (2022). Under various technical assumptions on the data distribution, such as strong log-concavity, there are Wasserstein convergence results available (e.g., Gao et al. (2025), Tang and Zhao (2024)), and it is possible to bound $\mathcal{A}\mathcal{W}_{2}$ by $\mathcal{W}_{2}$ up to some explicit error terms; see Blanchet et al. (2024). However, with this conventional approach, obtaining bounds for conditional distributions requires imposing strong log-concavity on every conditional distribution of $\mathbb{P}$. We circumvent these assumptions and replace them with certain disspative conditions of the networks, which are attained by carefully designing the network structure (i.e., Assumptions 3).
@@
    Taking $T=1$ in (2.31) or (2.35), we obtain from (2.36) a Wasserstein bound for static data generation:
    $\mathcal{W}_{2}^{2}(\mathbb{P},\mathbb{Q}^{\mathcal{T}})\leq C(\mathcal{T}^{\frac{5}{2}}\varepsilon_{\rm score}+\mathcal{T}^{2}e^{-c\mathcal{T}}\varepsilon_{\rm score}^{-1}),$
    which is interesting in its own right. This is because when $T=1$, Assumption 2 is reduced to the usual Lipschitz continuity of score functions, and, again, we do not rely on any structural assumptions on $\mathbb{P}$, such as strongly log concavity. This is achieved by bounding Wasserstein metric by total variation, up to some tail error (Blanchet et al. (2024), Proposition 20, and Villani (2008), Theorem 6.15), and making use of the uniform tail estimates of SDEs with dissipative coefficients. The price to pay is that the noise approximation error term $\mathcal{T}^{2}e^{-c\mathcal{T}}\varepsilon_{\rm score}^{-1}$ explodes when $\varepsilon_{\rm score}\to 0$.
@@
    It is evident that, as outlined in Algorithm 1, $\{y_{(n)}^{1}\}_{n=1}^{N}$ is sampled from the distribution $\mathbb{Q}_{1}^{\mathcal{T}}$. Furthermore, given the sequence $y_{(n)}^{1:t}$, the next sample $y_{(n)}^{t+1}$ is drawn from the distribution $\mathbb{Q}^{\mathcal{T}}_{y_{(n)}^{1:t}}$. By definition, $\{y_{(n)}^{1},y_{(n)}^{2},\cdots,y_{(n)}^{T}\}_{n=1}^{N}$ is one sample from the distribution $\mathbb{Q}^{\mathcal{T}}$, thereby establishing that $\mathbb{Q}^{\mathcal{T}}$ is the output distribution of Algorithm 1. Consequently, Theorem 2.3 provides $\mathcal{A}\mathcal{W}_{2}$ bounds between the output distribution of Algorithm 1 and the data distribution.
  • (4)
    Due to the construction of $\mathbb{Q}^{\mathcal{T}}$ (as defined in (2.34)), for a random variable $Y^{1:T}\sim\mathbb{Q}^{\mathcal{T}}$, $\mathbb{Q}^{\mathcal{T}}_{y^{1:t}}$ is the conditional distribution of $Y^{t+1}$ given $Y^{1:t}=y^{1:t}$. Consequently, it can be sampled by selecting the score matching function $s^{t+1}_{\theta}$. In essence, Algorithm 1 can also be utilized for conditional sampling, provided that $y^{1:t}$ is included as input, and the sampling process starts with $Z_{(n)}^{t+1}$.
@@
Consequently, for any $\varepsilon,p>0$, if $\mathcal{T}$ is sufficiently large such that $\alpha(\mathcal{T})<\varepsilon p/(2C)$, and $\varepsilon_{\rm score}<\frac{\varepsilon p\mathcal{T}^{-\frac{5}{2}}}{2C}$, then with probability at least 1p1-p, we have
@@
provided that the prior $X^{1:t}$ is randomly sampled from $\mathbb{P}_{1:t}$.